blake travels in a boat 200 miles downstream in the same time that it takes to go 120 miles upstream. If the speed of the stream is 5 mph, find the speed of the boat in still water
Answer by checkley77(12844)

200/(X+5)=120/(X-5)
120(X+5)=200(X-5)
120X+600=200X-1,000
120X-200X=-1,000-600
-80X=-1,600
X=-1,600/-80
X=20 MPH. IS THE SPEED OF THE BOAT IN STILL WATER.
PROOF:
200/(20+5)=120/(20-5)
200/25=120/15
8=8